OpenStudy (anonymous):

Simplify: 3(n+1)!/5n!

OpenStudy (anonymous):

3/5 n+1

OpenStudy (anonymous):

(n+1)! = (n+1) * n! (3/5)(n+1) (3n+3)/5

OpenStudy (saifoo.khan):

Imran cant be wrong!

OpenStudy (anonymous):

Could you include steps for how you got that? I actually wanna know how to do it

OpenStudy (anonymous):

it's same answer

OpenStudy (anonymous):

(n+1)(n)(n-1) --------------- n (n-1) every thing cancel beside n+1

jimthompson5910 (jim_thompson5910):

Remember, n! by definition means n! = n*(n-1)*(n-2)*(n-3)...(3)(2)(1) So (n+1)! = (n+1)(n)(n-1)(n-2)(n-3)...(3)(2)(1) Notice how n! and (n+1)! have the terms n*(n-1)*(n-2)*(n-3)...(3)(2)(1) in common. So when you divide (n+1)! by n!, these terms will cancel out. This will leave you with n+1 So in math notation, I'm saying that \[\frac{(n+1)!}{n!}=n+1\] This then means that \[\frac{3(n+1)!}{5n!}=\frac{3(n+1)}{5}=\frac{3n+3}{5}\]

OpenStudy (anonymous):

Ohh okay I think i get it. It cancels, cool thanks guys!